OpenStudy (anonymous):

The product of two consecutive negative integers is 1122. What are the numbers? The product of two consecutive negative integers is 1122. What are the numbers? @Mathematics

OpenStudy (lgbasallote):

let x be the first integer x + 1 = second integer the product of the two is 1122 so x(x+1) = 1122 x^2 + x = 1122 x^2 + x - 1122 = 0 use quadratic formula to solve for x